### v2

# K3 surfaces associated to a cubic fourfold

Abstract.
[…]
1. Introduction
Let $X$ be a complex smooth cubic fourfold in $\mathbf{P}^{5}$. Let $A(X)=H^{4}(X,\mathbf{Z})\cap H^{2,2}(X)$ be the lattice of algebraic cycles. The fourfold $X$ is special if the lattice $A(X)$ contains a class $v$ that is not homologous to $h^{2}$, where $h$ is the class of a hyperplane section. A fourfold $X$ is very general if ${\operatorname{rank}}A(X)=1$. For $X$ special let $K_{d}=<h^{2},v>$, where $d$ is the discriminant of $<h^{2},v>$.
A polarized K3 surface $(S,L)$ of degree $d$ and genus $g$, with $L^{2}=2g-2$ and $2g-2=d$, is said to be associated to a cubic fourfold $X\in\mathcal{C}_{d}$ if there is an isomorphism of Hodge structures
(1.1) $K^{\perp}_{d}\simeq H^{2}(S,\mathbf{Z})_{prim}(-1).$
Let $\mathcal{C}$ be the moduli space of smooth projective cubic fourfolds and let $\mathcal{C}_{d}\subset\mathcal{C}$ be the Hassett divisor parametrizing special cubic fourfolds with a labelling of discriminant d, i.e. a positive defined rank-two primitive sublattice $K_{d}=<h^{2},v>\subset A(X)$. B. Hassett in [Hass] proved that $X\in\mathcal{C}_{d}$ has an associated polarized K3 surface of degree $d$ if and only if $d$ satisfies the following numerical condition
(*) $d>6$ and $d$ is not divisible by 4,9 or a prime $p\equiv 2(3)$.
More precisely: for $d$ satisfying (*) $\mathcal{C}_{d}$ is birational to $\mathcal{N}_{d}$ if $d\equiv 2(6)$, where $\mathcal{N}_{d}$ is the moduli space of polarized K3 surfaces of degree d. Otherwise $\mathcal{C}_{d}$ is birational to a quotient of $\mathcal{N}_{d}$ by an involution, see [Hass, Cor.25].
The results of several Authors (Hassett,Kuznetsov, Addington-Thomas) suggest that a cubic fourfold is rational if and only if it has an associated K3 surface. Kuznetsov conjectured that X has an associated K3 surface if and only if there exists a semi-orthogonal decomposition of the derived category $D^{b}(X)$ of bounded complexes of coherent sheaves
(1.2) $D^{b}(X)=<\mathcal{A}_{X},\mathcal{O}_{X}(1),\mathcal{O}_{X}(2)>$
such that $\mathcal{A}_{X},$ is equivalent to the category $D^{b}(S)$,where $S$ is K3 surface.The following result proved in [BLMNPS, Cor.29.7], shows that the two conditions are in fact equivalent
Theorem 1.3.
Let $X$ be cubic fourfold. Then $X$ has a Hodge-theoretically associated K3 surface if and only if there exists a smooth projective K3 surface $S$ and an equivalence $\mathcal{A}_{X}\simeq D^{b}(S)$
These results leaded to the following Conjecture
Conjecture 1.4.
A smooth cubic fourfold $X$ is rational if and only if $X\in\mathcal{C}_{d}$ with $d$ satisfying the numerical condition in (*)
If $X$ is a very general cubic fourfold , i.e. $CH^{2}(X)\simeq\mathbf{Z}$, then $X$ has no associated K3 surface and hence is conjecturally irrational.
In the particular case where $d$ is an integer satisfying (*) which is of the form
$d=2(n^{2}+n+1)\ with\ n\in\mathbf{N}$
there is a rational map of degree 2
(1.5) $\phi:\mathcal{F}_{g}\to\mathcal{C}_{d},$
where $\mathcal{F}_{g}$ is the 19-dimensional moduli space of smooth polarized K3 surfaces $(S,H)$ of genus $g$ with $H^{2}=2g-2$ and $g=n^{2}+n+2$. The map $\phi$ sends $(S,H)$ to the class $|X|$ of a cubic fourfold $X\in\mathcal{C}_{d}$. The surface $S$ is associated to the cubic fourfold $X$ and there is an isomorphism
(1.6) $S^{[2]}\simeq F(X),$
where $S^{[2]}$ is the Hilbert scheme of two points on $S$.
Spelling out the numerical condition in (*) one finds that the above conjecture predicts rationality of all cubics in the Hasset divisors $\mathcal{C}_{d}$, for $d=14,26,38,42,62,74,78,86,98,....$
In [RS] it has been proved that a general $X\in\mathcal{C}_{d}$ with $d=14,26,38,42$ is rational. In all these cases the construction of a birational map $\mu:X\dashrightarrow W$, where $W$ is either $\mathbf{P}^{4}$ or a notable smooth rational fourfold, can be achieved by considering a surface $S_{d}\subset X$ admitting a four dimensional family of $(3e-1)$-secant curves of degree $e\geq 2$. These curves are parametrized by a rational variety with the property that trough a general point of $\mathbf{P}^{5}$ there passes a unique curve of the family. The inverse $\mu^{-1}:W\dashrightarrow X$ is given by a linear system of divisors in $|\mathcal{O}_{W}(e\cdot i(W)-1)|$, where $i(W)$ is the index of $W$, having points of multiplicity $e$ along an irreducible surface $U\subset W$, which is a birational incarnation of a K3 surface $S$ associated to $X$.
The results by Hassett and Kuznetsov have been extended by Bülles in [Bull] showing that if $X\in\mathcal{C}_{d}$, where $d$ satisfies the following numerical condition
$(**)\exists f,g\in\mathbf{Z}\ with\ g|2n^{2}+2n+2\ ,\ n\in\mathbf{N}\ and\ d=f^{2}g$
then there is a equivalence of categories
(1.7) $\mathcal{A}_{X}\simeq D^{b}(S,\alpha),$
with $S$ a K3 surface and $\alpha$ a Brauer class in $\operatorname{Br}(S)$. Clearly
$\bigcup_{d\in(*)}\neq\bigcup_{d\in(**)},$
For instance, if the fourfold $X$ contains a plane, and hence it belongs to $\mathcal{C}_{8}$ , then there is a K3 surface associated to $X$ in the sense of 1.7, but , for a general $X\in\mathcal{C}_{8}$, not in the Hodge theoretically sense.
Let $X$ be a special cubic fourfold and let $A(X)$ be the lattice of algebraic cohomology, i.e.
$A(X)=H^{4}(X,\mathbf{Z})\cap H^{2,2}(X)$
If ${\operatorname{rank}}A(X)\geq 12$ then $X$ has an associated K3 surface in the Hodge theoretically sense, see Remark. 2.11.
A third way in which a K3 surface can be possibly associated to a cubic fourfold is via the transcendental motives. Let $\mathcal{M}_{rat}(\mathbf{C})$ be the (covariant) category of Chow motives. The motive $h(X)$ of a cubic fourfold $X$ has a reduced Chow Künneth decomposition
(1.8) $h(X)=\mathbf{1}\oplus\mathbf{L}\oplus(\mathbf{L}^{2})^{\oplus\rho(X)}\oplus t(X)\oplus\mathbf{L}^{3}\oplus\mathbf{L}^{4},$
where $\mathbf{L}$ is the Lefschetz motive, $\rho(X)={\operatorname{rank}}A^{2}(X)$, with $A^{2}(X)=CH^{2}(X)\otimes\mathbf{Q}$ and $t(X)$ is the transcendental motive, see [BP]. Then
$H^{*}(t(X))=H^{4}_{tr}(X,\mathbf{Q})=T(X)_{\mathbf{Q}},$
where $T(X)\subset H^{4}(X.\mathbf{Z})$ is the transcendental lattice and
$A^{i}(t(X))=0\ for\ i\neq 3\ ;\ A^{3}(t(X))=A_{1}(X)_{hom}=A_{1}(X)_{alg},$
Similarly if $S$ is a K3 surface then its motive $h(S)$ has a reduced Chow-Künneth decomposition as follows
$h(S)=\mathbf{1}\oplus\mathbf{L}^{\oplus\rho(S)}\oplus t_{2}(S)\oplus\mathbf{L}^{2}$
where $\rho(S)$ is the rank of the Neron-Severi $N(SX)$, see [KMP].
Definition 1.9.
The motive $h(X)$ of a special cubic fourfold $X$ is associated to the motive of a K3 surface $S$ if there is an isomorphism of motives in the (covariant) category $\mathcal{M}_{rat}(\mathbf{C})$
(1.10) $t(X)\simeq t_{2}(S)(1),$
where $t_{2}(S)(1)=t_{2}(S)\otimes\mathbf{L}$.
The equivalence of categories in 1.7 implies that there is an isomorphism of transcendental motives, as in 1.10, see [Bull]. For a very general cubic fourfold no such isomorphism can exists, see [BP, Prop.3.4].
The existence of an isomorphism as in 1.10 implies that the motive $h(X)$ is finite-dimensional, in the sense of [Kim], if and only if $h(S)$ is such.
Some families of cubic fourfolds with a finite-dimensional motive have been described in [Lat 1] and [Lat 3]. In all these cases the motive $h(X)$ is of abelian type, i.e. lies in subcategory $\mathcal{M}^{Ab}_{rat}(\mathbf{C})$ generated by the motives of abelian varieties.
If a cubic fourfold $X$ has a finite group $G$ of symplectic automorphisms whose order order is $\neq 1,2$, then there is a unique K3 surface $S$ associated to $X$ in the sense of 1.2, i.e. $\mathcal{A}_{X}\simeq D^{b}(S)$, see [Ou, Thm.8.4]. Therefore $X$ is conjecturally rational and there is an isomorphism of motives as in 1.10.
In this note we consider families of cubic fourfolds admitting either an involution or a non-symplectic automorphism of prime order. Note that a non-symplectic prime order automorphism can have order 2 or 3.
In Section 2 we show that for every cubic fourfold $X$ with an involution the motive $h(X)$ is associated to the motive of a K3 surface $S$ in the sense of 1.10. For two distinct families of cubic fourfolds $X$ with an involution, there is no K3 surface Hodge theoretically associated to $X$, while there is one family consisting of cubic fourfolds with an involution that have an associated K3 surface, are rational and their motives are of abelian type.
In Sect 3 we deal with cubic fourfolds $X$ admitting a non -symplectic automorphism of order 3 and prove that there are 2 cases where $X$ has an associated K3 surface and is rational and 2 cases where $X$ is conjecturally irrational.
In Sect. 4 we describe families of special cubic fourfolds whose associated K3 surfaces are singular, i.e. their Picard rank is 20. These special cubic fourfolds have amotive of abelian type.
In Sect.5 we restrict to the case of cubic fourfolds $X$ such that $F(X)\simeq S^{2]}$, with $S$ a K3 surface, and show that the motive $h(X)$ is associated to the motive of $S$ in the sense of 1.10. We also show that a symplectic automorphism $\sigma$ on $S$ induces a symplectic automorphism $\tau$ on $X$ and give examples, in the case $\sigma$ has order 3, of K3 surfaces $S$ whose motive is associated to different cubic fourfolds $X_{1}\in\mathcal{C}_{d_{1}}$ and $X_{2}\in\mathcal{C}_{d_{2}}$ with $d_{1}\neq d_{2}$.
2. Cubic fourfold with an involution
The group of regular automorphisms of a cubic fourfold $X\subset\mathbf{P}^{5}$ is finite and any automorphism $f$ of $X$ is induced from a linear automorphism of the ambient projective space. Therefore the induced automorphism $f^{*}$ on $H^{4}(X,\mathbf{Z})$ preserves $h^{2}$, where $h$ is the class of a hyperplane section. The Abel-Jacobi map
$p_{*}q^{*}:H^{4}(X,\mathbf{Z})\simeq H^{2}(F(X),\mathbf{Z}),$
where $P_{F}\subset F(X)\times X$ is the universal family and $p,q$ are the projections to $F(X)$ and $X$, respectively, induces an isomorphism of Hodge structures
$H^{3,1}(X)\simeq H^{2,0}(F(X)).$
The automorphism $f$ is symplectic if $f^{*}$ acts as the identity on $H^{2,0}(F(X)$, that is equivalent to $f^{*}$ acting as the identity on $H^{3,1}(X)\simeq\mathbf{C}\omega$. If $f$ is non-symplectic and has order $n$ then
$f^{*}(\omega)=\zeta\omega,$
with $\zeta$ a primitive $n^{th}$-root of 1. Since $f$ has finite order we can assume that it acts diagonally, that is
$f(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})=(a_{0}x_{0},a_{1}x_{1},a_{2}x_{2},a_{3}x_{3},a_{4}x_{4},a_{5}x_{5}).$
Let’s denote by
$\det(f)=\prod_{i}a_{i}$
Let $F(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})=0$ be the equation of $X\subset\mathbf{P}^{5}$. If $X$ is invariant under the induced action of the automorphism $f$ there is a constant $\lambda_{f}\in\mathbf{C}^{*}$ such that
(2.1) $f^{*}(F)=\lambda_{f}F$
The action of $f$ is symplectic on $F(X)$ if and only if
$\det(f)/\lambda^{2}_{f}=1,$
see [BCS 1, Lemma 6.2].
There are three types of involutions on a cubic fourfold $X\subset\mathbf{P}^{5}$. Two of them are non-symplectic and one is symplectic, see [Marq]. An involution $\phi$ is symplectic if the induced automorphism on $H^{4}(X,\mathbf{Z})$ acts trivially on $H^{3,1}(X)$, it is non-symplectic if it acts as -1. Applying a linear change of coordinates we can assume that the involutions are the following ones
$\phi_{1}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to[x_{0},x_{1},x_{2},x_{3},x_{4},-x_{5}];$
$\phi_{2}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to[x_{0},x_{1},x_{2},x_{3},-x_{4},-x_{5}];$
$\phi_{3}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to[x_{0},x_{1},x_{2},-x_{3},-x_{4},-x_{5}].$
The involution $\phi_{2}$ is symplectic, while $\phi_{1}$ and $\phi_{3}$ are non-symplectic.
The family $\mathcal{F}_{1}$ of cubic fourfolds which are invariant under the involution $\phi_{1}$ has dimension 14 and coincides with the family of cubic fourfolds with an Eckardt point. If a smooth cubic 4-fold $X\subset\mathbf{P}^{5}$ contains an Eckardt point $p$ then we can choose coordinates $[x_{0},\cdots,x_{5}]$ in $\mathbf{P}^{5}$ such that $p=(0,0,0,0,0,1)$ and $X$ is defined by an equation
(2.2) $f(x_{0},x_{1},x_{2},x_{3},x_{4})+l(x_{0},x_{1},x_{2},x_{3},x_{4})x^{2}_{5}=0$
with $f$ of degree 3 and $l$ of degree 1,see [LZP,Lemmas 1.6 and 1.8].
Proposition 2.3.
Let $X$ be a cubic fourfold with an Eckardt point..There is a K3 surface $S$ with Picard rank 6 such that $t_{2}(S)(1)\simeq t(X)$. The K3 surface $S$ is not associated to $X$ in the Hodge theoretical sense.
Proof.
The equation $f(x_{0},x_{1},x_{2},x_{3},x_{4})=0$ in 2.2 defines a smooth cubic 3-fold $Y\subset\mathbf{P}^{4}$, with $\mathbf{P}^{4}:(x_{5}=0)$. Let $H\subset\mathbf{P}^{4}$ be the hyperplane defined by $l(x_{0},x_{1},x_{2},x_{3},x_{4})=0$. The intersection $S=Y\cap H$ is a cubic surface. The 27 planes $\mathbf{P}^{2}_{<l_{i},p>}$, where $(l_{1}\cdots,l_{27})$ are the lines on $S$, are contained in $X$. The lattice of the primitive algebraic cohomology $A(X)_{prim}$ is isomorphic to $E_{6}(2)$ and is spanned by the classes $[P_{i}]-[P_{j}]$,with $P_{i},P_{j}$ planes trough $p$, see [Marq,Thm.3.4]. There is a class $[F_{0}]\in A(X)$ such that $[F_{0}]\cdot[F_{0}]=7$ and $[F_{0}]\cdot h^{2}=3$, see [LZP, Lemma 2.4]. Therefore the rank two lattice $<h^{2},]F_{0}]>$ has discriminant 12. Since $X$ contains a plane we get $X\in\mathcal{C}_{8}\cap\mathcal{C}_{12}$.
Let $m$ be a line in the cubic $S=Y\cap H$ and let $P=\mathbf{P}^{2}_{<m,p>}\subset X$ be the plane generated by $m$ and the Eckardt poin $p$. Projecting $X$ from $P$ to a plane $P^{\prime}$ which is complementary to $P$ in $\mathbf{P}^{4}:(x_{5}=0)$ gives a quadric bundle
(2.4) $\pi:\tilde{X}\to P^{\prime}=\mathbf{P}^{2}$
whose discriminant locus is a degree 6 curve $D$. The curve $D$ consists of two irreducible components: a quintic curve $C$ and the line $L=H\cap P^{\prime}$. The quintic curve $C$ is the discriminant divisor of the conic bundle $\tilde{Y}\to P^{\prime}$, obtained by blowing-up the line $m$ on the cubic 3-fold $Y$. The fibre of $\pi$ over a singular point in $C\cap L$ is a couple of planes. Let $T\to\mathbf{P}^{2}$ be the double cover parametrizing the lines in the fibres of $\pi$. Let $F(\tilde{X}/\mathbf{P}^{2})$ be the relative Fano variety of lines in the fibration $\pi$ and let $\mathcal{L}\to F(\tilde{X}/\mathbf{P}^{2})$ be the universal line sitting in the diagram
[…]
In the Stein factorization of the composite map $\mathcal{L}\to\mathbf{P}^{2}$
(2.6) $\mathcal{L}\to T\to\mathbf{P}^{2}$
the map $\mathcal{L}\to T$ is a $\mathbf{P}^{1}$ bundle and $T\to\mathbf{P}^{2}$ is a double cover ramified along the reduced sextic $C\cup L$. Let $\tau:\tilde{\mathbf{P}}^{2}\to\mathbf{P}^{2}$ be the blow-up of $\mathbf{P}^{2}$ at the 5 nodes in $C\cap L$ and let $S\to\tilde{\mathbf{P}}^{2}$ be the double cover ramified along the strict transform of $D=C\cup L$. Then $S$ is a K3 surface and $\tau|_{S}:S\to T$ is a desingularization. Let
$f:S\to T\to\mathbf{P}^{2}$
and let $h=f^{*}[l]$ be the pull-back of the class of a line from $\mathbf{P}^{2}$. Since the surface $T$ has five ordinary double points coming from the intersection $C\cap L$ it follows that $\operatorname{Pic}S$ contains five additional classes corresponding to the exceptional divisors of $S\to T$. The line $L$ is in the branch locus of $f$ and hence the surface $S$ contains another rational curve, namely the inverse image $L^{\prime}$ of the line $L$. Therefore the K3 surface $S$ contains 6 irreducible (-2)-curves whose classes $l^{\prime},e_{1},\cdots,e_{5}$ span a rank 6 lattice $M$ which admits a unique primitive embedding in $\operatorname{Pic}S$. The cubic fourfold $X$ is not associated to the K3 surface $S$ in the Hodge theoretically sense see [Laz , Rk.3.9].
Since $X$ contains a plane there is an isomorphism of motives
$t_{2}(S)(1)\simeq t(X),$
see [Bull, Ex 3.3]. ∎
Proposition 2.7.
Let $\mathcal{F}_{2}$ be the family of cubic fourfolds that are invariant under the involution $\phi_{2}$. If $X\in\mathcal{F}_{2}$ there is a K3 surface $S$, double cover of a cubic surface branched along a degree 6 curve, such that
$t(X)\simeq t_{2}(S)(1)$
The cubic $X$ belongs to $\mathcal{C}_{12}$, contains no plane and has no Hodge theoretically associated K3 surface,
Proof.
A cubic $X$ invariant under $\phi_{2}$ has an equation of the form
(2.8) $f(x_{0},x_{1},x_{2},x_{3})+x^{2}_{4}l_{1}+x^{2}_{5}l_{2}+x_{4}x_{5}l_{3}=0,$
where $f$ is homogeneous of degree 3 and $l_{i}$ are homogeneous of degree 1 in $x_{0},x_{1},x_{2},x_{3}$. The involution $\phi_{2}$ is symplectic. The locus of fixed points of $\phi_{2}$ on $\mathbf{P}^{5}$ is the disjoint union of a $\mathbf{P}^{3}$ defined by $x_{4}=x_{5}=0$ and the line $r$ joining the base points $P_{4}$ and $P_{5}$. The line $r$ is contained in $X$ and the fixed locus on $P^{3}$ is the cubic surface $W:f(x_{0},x_{1},x_{2},x_{3})=0$.The fixed locus of $\phi_{2}$ on $F(X)$ consists of the line $r$, the 27 lines on a cubic surface $W$ and a K3 surface $S$, see [Cam]. The surface $S$ parametrizes the lines joining a fixed point $Q_{1}$ on $\mathbf{P}^{3}$ and a point $Q_{2}$ on $r$. The surface $S$ is a double cover of the cubic surface $W$ branched along a degree 6 curve , see [BP, rk.4.4].
The cubic $X$ contains no plane and has no associated K3 surface, see [Marq, Lemma 4.1 and Lemma 6.3]. The lattice $A(X)_{prim}$ has rank 8 and is generated by the classe $[T_{i}]-h^{2}$, where $T_{i}$ are cubic scrolls. Therefore $X\in\mathcal{C}_{12}$. There is an isomorphism
$t_{2}(S)(1)\simeq t(X)$
as in 1.10, see [BP, Prop.4.6]. ∎
The family $\mathcal{F}_{3}$ of cubic fourfolds with the anti-symplectic involution $\phi_{3}$ has dimension 10. Every $X\in\mathcal{F}_{3}$ has an equation of the form
$f(x_{0},x_{1},x_{2})+x_{0}q_{0}(x_{3},x_{4},x_{5})+x_{1}q_{1}(x_{3},x_{4},x_{5})+x_{2}q_{2}(x_{3},x_{4},x_{5})=0,$
with $f$ homogeneous of degree 3 and $q_{0},q_{1},q_{2}$ homogeneous of degree 2. In [Marq, Thm. 6.4] it is proved that $\mathcal{F}_{3}\subset\mathcal{I}$ where
$\mathcal{I}=\bigcap_{d>6,d\equiv 0,2(6)}\mathcal{C}_{d}$
is the intersection of all non-empty Hasset divisors. In [YY] it is proved that $\mathcal{I}$ is non- empty, contains the Fermat cubic, consists of rational cubic fourfolds and
$13\leq\dim\mathcal{I}\leq 16$
Lemma 2.9.
Every cubic fourfold $X\in\mathcal{I}$ has an associated K3 surface with a motive of abelian type.
Proof.
By the results in [ABP, Appendix A] one can choose suitable integers $d_{1},\cdots,d_{19}$, with $d_{1}=14$, such that every $X\in\bigcap_{1\leq i\leq 19}\mathcal{C}_{d_{i}}$ has an associated K3 surface whose Picard rank equals 19 and hence has a motive of abelian type, see [Ped]. Therefore the motive $h(X)$ is of abelian type. If $X\in\mathcal{I}$ then also $X\in\bigcap_{1\leq i\leq 19}\mathcal{C}_{d_{i}}$. ∎
Proposition 2.10.
Every cubic fourfold $X\in F_{3}$ is rational, has an associated K3 surface $S$ and there is an isomorphism of motives
$t_{2}(S)(1)\simeq t(X)$
where $t_{2}(S)$ and $t(X)$are of abelian type.The lattice $A(X)$ of the algebraic cohomology has rank 11.
Proof.
Since $F_{3}\subset\mathcal{I}$ every $X\in F_{3}$ is rational. By 2.9 the cubic $X$ has an associated K3 surface whose motive is of abelian type.
The plane $P:(x_{0}=x_{1}=x_{2}=0)$ is contained in $X$ and is invariant under $\phi_{3}$. By [Bull, Ex.3.3] There is an isnomorphism
$t_{2}(S)(1)\simeq t(X)$
with $S$ a K3 surface. The cubic $X$ is Pfaffian, see [Marq, Prop.6.18]. ∎
Remark 2.11.
If $X$ is a cubic fourfold such that ${\operatorname{rank}}A(X)\geq 12$ then the transcendental lattice $T(X)$ has a primitive embedding into the K3 lattice $(E_{8})^{2}\oplus U^{3}$,see [Laz ,Prop.2.9]. Therefore the maximal rank of $A(X)$ that can occur for a cubic not associated to a K3 surface, and hence conjecturally irrational, is 11. In [YY,Cor 8.18] it is proved that there are infinitely many rank 11 lattices as the algebraic cohomology of smooth cubic fourfolds which do not admit a Hodge theoretically associated K3 surface. On the other hand the 10-dimensional family $\mathcal{F}_{3}$ consists of rational cubic fourfolds with an associated K3 surface and whose algebraic cohomology has rank 11.
3. Automorphisms of order 3
A smooth cubic fourfold $X\subset\mathbf{P}^{5}$ can have prime order automorphisms of orders $2,3,5,7,11$. In [GAL,Thm.3.8] there is a list of the automorphisms of $\mathbf{P}^{5}$ of order $2,3,5,7,11$ and a description of the families of cubic hypersurfaces preserved by their action. G.Ouchi in [Ou,Thm.8.4] proved that if a smooth cubic fourfold has a symplectic group of automorphisms $G$, whose order is different from 1 and 2, then ${\operatorname{rank}}S_{G}(X)\geq 12$, where
(3.1) $S_{G}(X)=(H^{4}(X,\mathbf{Z})^{G})^{\perp}.$
Moreover there exists a unique K3 surface $S$ such that $\mathcal{A}_{X}\simeq D^{b}(S)$, with $\mathcal{A}_{X}$ the Kuznetsov component in the derived category $D^{b}(X)$. Therefore $X$ has a Hodge theoretically associated K3 surface $S$, it is conjecturally rational and there is an isomorphisms of motives $t_{2}(S)(1)\simeq t(X)$.
In this section we consider the case of non-symplectic automorphisms of order 3 and show when $X$ has an associated K3 surface and is rational and when $X$ has no associated K3 surface.
In [GAL,Thm.3.8] there is a list of the automorphisms of $\mathbf{P}^{5}$ of order 3 and a description of the families of cubic hypersurfaces $X$ such that the automorphism acts non-symplectically on $X$.
Let
$\sigma_{1}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to:[x_{0},x_{1},x_{2},x_{3},x_{4},\zeta x_{5}];$
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1 and let $V_{1}$ be the family of cubic fourfolds invariant under $\sigma_{1}$. The automorphism $\sigma_{1}$ induces a non-symplectic automorphism on $F(X)$.The family $V_{1}$ has dimension 10.
Proposition 3.2.
. Let $X$ be a general element in $V_{1}$. The cubic fourfold $X$ has no Hodge theoretically associated K3 surface. The motive $h(X)$ is finite-dimension and is not associated to the motive of a K3 surface.The Fano variety $F(X)$ is not isomorphic to the Hilbert scheme of points $S^{[2]}$ of a K3 surface $S$.
Proof.
A cubic fourfold i $X\in V_{1}$ has an equation of the form
$F(x_{0},x_{1},x_{2},x_{3},x_{4})+x^{3}_{5}=0$
with $F$ homogeneous of degree 3. The action of $\sigma_{1}$ is non -symplectic because $\det\sigma_{1}=\zeta$ and $\sigma(F)=F$.The cubic fourfold $X$ is a triple cover of $\mathbf{P}^{4}$ ramified along the cubic 3-fold $Y:(F(x_{0},x_{1},x_{2},x_{3},x_{4})=0)$.The dimension of the family $V_{1}$ is 10, see [GAL]. All these fourfolds have a motive of abelian type, see [Lat 1].
[…]
Let $T$ be the invariant lattice of $\sigma_{1}$ in $H^{2}(F,\mathbf{Z})$ and let $S=T^{\perp}$. Then ${\operatorname{rank}}S=22$, and $T=<6>=<g,g>$, with $g$ the restriction to $F\subset\operatorname{Grass}(2,6)$ of the class defining the Plücker embedding, see [BCS 1 (6.3)]. Let $P_{F}\to F$ be the universal family and denote by $p$ and $q$ the projection to $F$ and $X$ respectively. Then the Abel-Jacobi map
(3.3) $p_{*}q^{*}:H^{4}(X,\mathbf{Z})\to H^{2}(F,\mathbf{Z})$
induces an isomorphism between the primitive cohomology $H^{4}(X,\mathbf{Z})_{prim}$ and $H^{2}(F,\mathbf{Z})_{prim}$. Here $H^{2}(F,\mathbf{Z})_{prim}=<g>^{\perp}$, $g=p_{*}q^{*}(h^{2})$, with $h$ the class of a hyperplane section and
(3.4) $CH^{1}(F)_{\mathbf{Q}}\simeq<g>\oplus p_{*}q^{*}(A^{2}(X)_{prim},$
where
$A^{2}(X)_{prim}=\{Z\in CH^{2}(X)_{\mathbf{Q}}/[Z]\in H^{4}(X,\mathbf{Q})_{prim}\}$
see [SV, 21.4].
For a general element $X\in V_{1}$ the group $NS(F(X))$ is equal to the invariant subgroup $T$ which has rank 1. From the isomorphisms in 3.4 we get
$A_{2}(X)_{prim}=0,$
Therefore the algebraic lattice $A(X)$ has rank 1 and the cubic fourfold $X$ has no associated K3 surface, in the Hodge theoretical l sense.
Suppose that there exists a K3 surface $S$ such that
$t(X)\simeq t_{2}(S)(1),$
Then
$H^{4}(t(X))=H^{4}(X,\mathbf{Q})_{tr}\simeq H^{4}(t_{2}(S)(1))=H^{2}(S,\mathbf{Q})_{tr}.$
Since the dimension of $H^{2}(S,\mathbf{Q})_{tr}$ is at most 21 we should also have $\dim H^{4}(X,\mathbf{Q})_{tr}\leq 21$, while $\dim H^{4}(X,\mathbf{Q})_{tr}=22$, because ${\operatorname{rank}}A^{2}(X)=1$.
If $X$ is a general element of $V_{1}$ the Fano variety $F(X)$ is not isomorphic to $S^{[2]}$, with $S$ a K3 surface, because $NS(F(X)$ equals the invariant lattice that has rank 1 while $NS(S^{[2]})$ has rank at least 2. ∎
Remark 3.5.
(1) R.Laterveer in [Lat 2, Cor.5.2] proved that $A_{0}(Z)_{hom}=0$, with $Z=F(X)/<\sigma_{1}>$ and $X\in V_{1}$.
(2)Since ${\operatorname{rank}}A(X)=1$ the family $V_{1}$ contains contains no plane. The family $V_{1}$ contains a Pfaffian cubic fourfold $Y$ (see [BCS 2,Prop.5.1]. Since $Y\in\mathcal{C}_{14}$ there is an isomorphism $F(Y)\simeq S^{[2]}$, where $S$ is a K3 surface and an isomorphism $t_{2}(S)(1)\simeq t(Y)$. The K3 surface $S$ has a motive of abelian type.
(3) The family $V_{1}$ contains the Klein cubic fourfold $X_{K}$ whose equation is of the form
$x^{2}_{0}x_{1}+x^{2}_{1}x_{2}+x^{2}_{2}x_{3}+x^{2}_{3}x_{4}+x^{2}_{4}x_{0}+x^{3}_{5}=0.$
It is a triple cover of $\mathbf{P}^{4}$ branched along the Klein cubic 3-fold. The fourfold $X_{K}$ admits also a symplectic automorphism of order 11 and therefore has an associated K3 surface.
Let
$\sigma_{2}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to:[x_{0},x_{1},x_{2},x_{3},\zeta x_{4},\zeta x_{5}];$
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1. Let $V_{2}$ be the family of cubic fourfolds invariant under $\sigma_{2}$. The family $V_{2}$ has diimension 4. Every cubic fourfold $X\subset\mathbf{P}^{5}$ is defined by an equation of the form
(3.6) $F(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})=f(x_{0},x_{1},x_{2},x_{3})+g(x_{4},x_{5})=0,$
with $f$ and $g$ homogeneous of degree 3 .The automorphism $\sigma_{2}$ acts non-symplectically on $F(X)$.
Proposition 3.7.
A general cubic $X\in V_{2}$ has an associated K3 surface $S$, belongs to $\mathcal{C}_{14}$ and is rational. The motives $h(X)$ and $h(S)$ are both of abelian type.
Proof.
The automorphism $\sigma_{2}$ acts non-symplectically because $\det\sigma_{2}=\zeta^{2}$ and $\sigma^{*}(F)=F$.
The fixed locus of $\sigma_{2}$ on $X$ is
$P:\{x_{0}=x_{1}=x_{2}=x_{3}=0\ ,\ G(x_{4},x_{5})=0\}$
consisting of 3 points $p_{1},p_{2},p_{3}$ and the cubic surface
$S:\{x_{4}=x_{5}=0\ ,\ f(x_{0},x_{1},x_{2},x_{3})=0\}.$
The 3 points $p_{i}$ lie on the line $x_{0}=x_{1}=x_{2}=x_{3}=0$ which is not contained in $X$. Let $l$ be a line joining a point $p_{i}$ and a point $Q\in S$. The line $l$ is invariant and is contained in $X$, because otherwise the third point $T$ of intersection in $l\cap X$ would be a fixed point of $\sigma$ while $T\neq p_{i},T\notin S$. Therefore on $F(X)$ the automorphism induced by $\sigma_{2}$ has three fixed surfaces $K_{1},K_{2},K_{3}$ each one isomorphic to the rational cubic $S$. Moreover each fixed line on $S$ determines a fixed point on $F(X)$ so we get 27 isolated fixed points.
Let $K_{1}=F(X)_{p_{1}}$ and $K_{2}=F(X)_{p_{2}}$ be the two cubic surfaces in $F(X)$ which are the locus of lines joining the point $p_{1}$ and the point $p_{2}$ with the points on $S$. There are 27 planes contained in $X$ and passing trough $p_{1}$ i.e. the planes $P^{(i)}_{1}=\mathbf{P}^{2}_{<l_{i},p_{1}>}$, with $(l_{1},\cdots,l_{27})$ the lines on $S$. Similarly there are 27 planes contained in $X$ and passing trough $p_{2}$, i.e. the planes $P^{(i)}_{2}=\mathbf{P}^{2}_{<l_{i},p_{2}>}$. The planes $P^{(i)}_{1}$ and $P^{(i)}_{2}$ meet along the line $l_{i}$ while two planes $P^{(i)}_{1}$ and $P^{(j}_{2}$, with $i\neq j$ , $i,j=1\cdots 27$, meet in one point $r_{ij}$ if $l_{i}\cap l_{j}\neq\emptyset$. Since a smooth cubic fourfold has at most 10 planes with one point intersection, see [DM, Cor.7.3] there are some of the planes $P^{(i)}_{1},P^{(j)}_{2}$ which are contained in $X$ and do not intersect. Every cubic fourfold $X$ containing two disjoint planes $P,P^{\prime}$ belongs to $\mathcal{C}_{8}\cap\mathcal{C}_{14}$, hence it is rational and has an associated K3 surface $S$, see [Hass ,1.2].
The invariant lattice $T_{\sigma_{2}}(X)\subset H^{2}(F(X),\mathbf{Z})$ has rank 13, see [BCS 1, Ex.6.5], and hence the isomorphism in 3.4 implies ${\operatorname{rank}}A(X)\geq 13$.
Let $X_{1},X_{2}$ be the cubics of dimension 3 and 1, respectively
$X_{1}:f(x_{0},x_{1},x_{2},x_{3})-y^{3}=0\ ,\ X_{1}\subset\mathbf{P}^{4}$
$X_{2}:g(x_{4},x_{5})+z^{3}=0\ ,\ X_{2}\subset\mathbf{P}^{2}$
The rational map $\mathbf{P}^{4}\times\mathbf{P}^{2}\dashrightarrow\mathbf{P}^{5}$
$\Phi:[x_{0},x_{1},x_{2},x_{3},y;x_{4},x_{5},z]\to[x_{0}/y,x_{1}/y,x_{2}/y,x_{3}/y,x_{4}/z,x_{5}/z]$
induces a dominant rational map $\Psi$ of degree 3 of $X_{1}\times X_{2}$ on the hypersurface in $\mathbf{P}^{5}:(z_{0},z_{1},z_{2},z_{3},z_{4},z_{5}),$ defined by the equation
$f(z_{0},z_{1},z_{2},z_{3})+g(z_{4},z_{5})=0$
which is isomorphic to $X$, see [CT, Prop.1.2] and [Lat 3,Prop.6].
The indeterminacy locus of $\Psi$ is
$Y=(X_{1}\cap(y=0)\times(X_{2}\cap(z=0).$
Blowing up $X_{1}\times X_{2}$ along $Y$ gives a finite surjective morphism $\overline{X_{1}\times X_{2})}\to X$, with $\overline{X_{1}\times X_{2}}=\operatorname{Bl}_{Y}(X_{1}\times X_{2})$. Therefore $h(X)$ is a direct summand of $h(\overline{X_{1}\times X_{2}})$ and
$h(\overline{X_{1}\times X_{2}})\simeq h(X_{1}\times X_{2})\oplus h(Y)(1).$
Since $(X_{1}\cap(y=0)$ is a cubic surface and $(X_{2}\cap(z=0)$ consists of 3 points the motive $h(Y)$ has no transcendental part. Therefore the transcendental motive $t(X)$ is a direct summand of $h(X_{1}\times X_{2})$ where
$h(X_{1}\times X_{2})=\sum_{0\leq i\leq 8}h_{i}(X_{1}\times X_{2})\ ;\ h_{i}(X_{1}\times X_{2})=\sum_{p+q=i}h_{p}(X_{1})\times h_{q}(X_{2}).$
The motive of the cubic 3-fold $X_{1}$ has a Chow-Künneth decomposition
$h(X_{1})\simeq\mathbf{1}\oplus\mathbf{L}\oplus N\oplus\mathbf{L}^{2}\oplus\mathbf{L}^{3},$
where $N=h_{1}(J)(1)$ with $J$ an abelian variety isogenous to the intermediate Jacobian $J^{2}(X_{1})$, see[GG]. The motive $h(X_{2})$ of the elliptic curve $X_{2}$ is isomorphic to $\mathbf{1}\oplus h_{1}(X_{2})\oplus\mathbf{L}$ and is of abelian type. Therefore the transcendental motive $t(X)$ is a direct summand of $(h_{1}(J)\otimes h_{1}(X_{2})(2)$ and is of abelian type. Since $X\in\mathbf{C}_{14}$, there is an isomorphism of motives
$t_{2}(S)(1)\simeq t(X)$
and therefore also the K3 surface $S$ has a motive of abelian type. ∎
Proposition 3.8.
Let
$\sigma_{3}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to:[x_{0},x_{1},x_{2},\zeta x_{3},\zeta x_{4},\zeta^{2}x_{5}];$
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1 and let $V_{3}$ be the family of cubic fourfolds with a non-symplectically action of the automorphism $\sigma_{3}$.The family $V_{3}$ has dimension 7. For a general element $X\in V_{3}$ the lattice $A(X)$ has rank 7 and it is generated by the classes of 27 cubic scrolls. The fourfold $X$ belongs to the Hasset divisor $\mathcal{C}_{12}$, contains no plane and has no Hodge theoretically associated K3 surface. There is an isomorphism of motives in $\mathcal{M}_{rat}(\mathbf{C})$
$t_{2}(S)(1)\simeq t(X)$
where $S$ is a K3 surface.
Proof.
Every $X\in V_{3}$ is defined by an equation of the form
$F(x_{o},x_{1},x_{2},x_{3},x_{4},x_{5})=f(x_{0},x_{1},x_{2})+g(x_{3},x_{4})+x^{3}_{5}+$
$+x_{3}x_{5}l_{1}(x_{0},x_{1},x_{2})+x_{4}x_{5}l_{2}(x_{0},x_{1},x_{2})=0,$
where $f,g$ are homogeneous of degree 3 and $l_{1},l_{2}$ are linear forms. The automorphism $\sigma_{3}$ is non- symplectic because $\det\sigma_{3}=\zeta$ and $\sigma_{3}(F)=F$.The family $V_{3}$ has dimension 7, see [GAL,Thm.3.8].
In [BG, Lemma 4.9 and Prop.4.11] it is proved that $X\notin\mathcal{C}_{d}$, for $d\equiv 2(6)$. Therefore $X$ contains no plane. The fourfold $X$ contains 27 families of cubic scrolls $\{[T_{i}],[T^{*}_{i}]\}$, $i=1,\cdots 27$, such that $A(X)$ is generated by $[T_{i}]$.Therefore $X\in\mathcal{C}_{12}$. Since $d=12$ satisfies the numerical condition in (**) there is an isomorphism of motives as in 1.9.
There is no primitive embedding of $T(X)(-1)$ in a K3 lattice, hence there is no associated K3 surface, see [BG,lemma 4.7].
Remark 3.9.
The family $V_{3}$ contains the cubic fourfold $Y$ defined by the equation
$Y:f(x_{0},x_{1},x_{2})+g(x_{3},x_{4})+x^{3}_{5}=0$
which is also invariant under the symplectic automorphism
$\tau:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to[x_{0},x_{1},x_{2},\zeta x_{3},\zeta x_{4},\zeta x_{5}].$
The cubic fourfold $Y$ is rational and has a motive of abelian type, see [CT] and [Lat 3]. The fixed locus $\sigma_{3}$ on $F(X)$, for a general $X\in V_{3}$, is parametrized by $C\times D$, were $C,D$ are the elliptic curves defined by
$C:f(x_{0},x_{1},x_{2})=0\ in\ \mathbf{P}^{2}:x_{3}=x_{4}=x_{5}=0,$
$D:g(x_{3},x_{4})+x^{5}=0\ in\ \mathbf{P}^{2}:x_{0}=x_{1}=x_{2}=0$
and coincides with the fixed locus for the action of $\tau$ on $F(Y)$. In [Kaw,Prop.3.2] it is proved that the quotient variety $F(Y)/<\tau>$ has a crepant resolution $\tilde{Y}\to F(Y)/<\tau>$, with $\tilde{Y}$ birational to the generalized Kummer variety $K^{2}(C\times D)$. The same argument shows that the motive of the quotient $F(X)/<\sigma_{3}>$ is a direct summand of $K^{2}(C\times D)$, hence it is of abelian type.
Let
$\sigma_{4}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to:[x_{0},x_{1},\zeta x_{2},\zeta x_{3},\zeta^{2}x_{4},\zeta^{2}x_{5}].$
be the automorphism of $\mathbf{P}^{5}$, where $\zeta$ is a primitive cubic root of 1. There are two families of cubic fourfolds which are invariant under the action of $\sigma_{4}$, a first one where $\sigma_{4}$ acts non-symplectically and a second one where $\sigma_{4}$ acts symplectically, see [GAL,Thm. 3.8]. Let $V_{4}$ be the family of cubic fourfolds with a non-symplectical action of $\sigma_{4}$. Every $X\in X$ has an equation of the form
$F(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})=x_{2}(L_{2}(x_{0},x_{1})+x_{3}M_{2}(x_{0},x_{1})+x^{2}_{4}L_{1}(x_{0},x_{1})+$
$+x_{4}x_{5}M_{1}(x_{0},x_{1})+x^{2}_{5}N_{1}(x_{0},x_{1})+x_{4}N_{2}(x_{2},x_{3})+x_{5}P_{2}(x_{2},x_{3})=0$
where $L_{2},M_{2},N_{2},P_{2}$ are homogeneous polynomial of degree 2 and $L_{1},M_{1},N_{1}$ are linear factors. The automorphism $\sigma_{4}$ acts non-symplectically on $F(X)$ because $\det\sigma_{4}=1$ and $\sigma^{*}_{4}(F)=\zeta$.
Proposition 3.10.
The family $V_{4}$ has dimension 6. A general fourfold $X\in V_{4}$ contains nine disjoint planes $F_{1},\cdots,F_{9}$ and $A(X)$ has a basis given by $\{h^{2},[F_{1}],\cdots,[F_{8}]\}$. Therefore $X\in\mathcal{C}_{14}$, is rational and has an associated K3 surface.
Proof.
A plane in $\mathbf{P}^{5}$ that is invariant for the action of $\sigma_{4}$ has an equation of the form
$\{ax_{0}=bx_{1},cx_{2}=dx_{3},ex_{4}=fx_{5}\}.$
There are exactly nine disjoint planes $F_{1},\cdots,F_{9}$ that are invariant under $\sigma_{4}$ and are contained in $X$. Moreover
$h^{2}=(1/3)\sum_{1\leq i\leq 9}[F_{i}],$
see [BG, Prop. 4.13]. ∎
4. Singular K3 surfaces
In this section we describe families of special cubic fourfolds $X$ with an associated K3 surface $S$ whose Picard rank equals to 20, i.e. a singular K3 surface. These K3 surfaces have a motive of abelian type, see [Ped]. Since the motive $h(X)$ is associated to the motive of $S$, in the sense of 1.9, also $h(X)$ is of abelian type.
The following result appears in [ABP ,Thm.1.1] : let $3\leq n\leq 20$ and let $d_{k}\geq 8,d_{k}\equiv 0,2[6]$, where
[…]
with $p_{i}$ a prime number. Then
(4.1) ${\bigcap_{k=1}^{n}}\mathcal{C}_{d_{k}}\neq\emptyset,$
Note that the first two discriminants $d_{1}$ and $d_{2}$ may be chosen completely arbitrarily. A very general cubic fourfold $X$ has $CH_{2}(X)=\mathbb{Z}$, and each time that we take the intersection with a divisor $\mathcal{C}_{d_{i}}$ we know that we are adding an algebraic cycle inside $A(X)$. Therefore the generic cubic fourfold in ${\bigcap_{k=1}^{n}}\mathbf{C}_{d_{k}}\neq\emptyset$ has ${\operatorname{rank}}(CH_{2}(X))=n+1$. If one of the discriminants $d_{i}$ defines a divisor of cubics with associated K3 surface, it means that all the cubic fourfolds inside the intersection 4.1 have associated K3 surfaces and, by the definition itself, the generic associated K3 surface has Picard rank equal to $n$.
Let ${\bigcap_{k=1}^{20}}\mathbf{C}_{d_{k}}\neq\emptyset$, with at least one of the divisor $\mathcal{C}_{d_{k}}$ parametrizing cubic fourfolds with an associate K3 surface. For every $X\in\displaystyle{\bigcap_{k=1}^{20}}\mathbf{C}_{d_{k}}$ the lattice $A(X)$ of algebraic cycles has rank 21, hence the primitive cohomology of the associated K3 surface has rank 19 and the Neron-Severi group has rank 20.
The singular K3 surface $S$ associated to a cubic fourfold $X\in\displaystyle{\bigcap_{k=1}^{20}}\mathbf{C}_{d_{k}}$ admits a Shioda-Inose structure. In particular there is a birational map from $S$ to a Kummer surface, which is the quotient of an Abelian variety by an involution.
Therefore the motive $h(X)$ is associated to $h(S)$ and are both of abelian type. The Fano variety $F(X)$ is birational to the Hilbert scheme of length 2 subschemes of a Kummer surface.
A different way to construct cubic fourfolds with associated singular K3 surfaces is to consider a finite group $G$ of symplectic automorphisms acting on $X$ such that the covariant lattice $S_{G}(X)$ has maximal rank. The moduli space parametrizing cubic fourfolds with a symplectic group $G$ has dimension $20-{\operatorname{rank}}S_{G}$. The case ${\operatorname{rank}}S_{G}=20$ corresponds to cubic fourfolds $X$ such that the lattice of algebraic cycles $A(X)=<S_{G},h^{2}>$ , with $h$ the class of hyperplane section, has rank 21. All these cubics have an associated K3 surface $S$ with Picard rank equal to 20. Therefore all these cubic fourfolds have a motive of abelian type.
In [LZ,Thm.1.8] a list is given of all the groups $G$ which are the symplectic automorphism group of a cubic fourfold $X$ with ${\operatorname{rank}}S_{G}(X)=20$. There are six of such groups and 8 associated cubic fourfolds. The list contains the Fermat cubic $X_{F}$, the Klein cubic fourfold $X_{K}$ and the Clebsch diagonal cubic $X_{C}$. The Fermat cubic $X_{F}$ is the only smooth cubic fourfold with a symplectic automorphism of order 9. It is rational and is contained in the intersection of all Hasset divisors $\mathcal{C}_{d}$, with $d$ satisfying (*). The Klein cubic fourfold $X_{K}$ is the only cubic with a symplectic automorphism of order 11. It is contained in all $\mathcal{C}_{3d}$ with $d$ even and $d>4$.
The diagonal cubic $X_{C}$ has equation
$F(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})=x^{3}_{0}+x^{3}_{1}+x^{3}_{2}+x^{3}_{3}+x^{3}_{4}+x^{3}_{5}-(x_{0}+x_{1}+x_{2}+x_{3}+x_{4}+x_{5})^{3}=0.$
and has a symplectic automorphism of order 7.It is i also invariant with respect to a anti-symplectic involution, see [LZ,Cor.6.9]. Therefore $X_{C}$ contains an Eckardt points and belongs to $\mathcal{C}_{8}\cap\mathcal{C}_{12}$.
5. K3 surfaces with a symplectic automorphism
In the previous sections we described several families of cubic fourfolds $X$, with a finite group of automorphisms, such that the motive $h(X)$ is associated to the motive of a K3 surface $S$. Here we consider the case of K3 surface $S$ such that $S^{[2]}\simeq F(X)$ and show that $t_{2}(S)(1)\simeq t(X)$. If $S$ is equipped with a symplectic automorphism of prime order $p$ then also $X$ has a symplectic automorphism of order $p$. In the case $p=3$ we give examples of K3 surfaces $S$ whose motive is associated to different cubic fourfolds $X_{1}\in\mathcal{C}_{d_{1}}$ and $X_{2}\in\mathcal{C}_{d_{2}}$ with $d_{1}\neq d_{2}$.
Let $\mathcal{F}_{g}$ be the moduli space of smooth polarized K3 surfaces $(S,L)$ of genus $g$ and degree $d$, with $L^{2}=d=2g-2$ and $g=n^{2}+n+2$. There is a rational map of degree 2
(5.1) $\phi:\mathcal{F}_{g}\dashrightarrow\mathcal{C}_{d},$
where $d=2(n^{2}+n+1)$, sending $(S,L)$ to the class $|X|$ of a cubic fourfold $X\in\mathcal{C}_{d}$. There is an isomorphism
(5.2) $S^{[2]}\simeq F(X).$
Lemma 5.3.
Let $X$ be a cubic fourfold and let $F=F(X)$ be the Fano variety of lines. Suppose that $F\simeq S^{[2]}$, with $S$ a K3 surface. Then there is an isomorphism of motives $t_{2}(S)(1)\simeq t(X)$ in $\mathcal{M}_{rat}(\mathbf{C})$.
Proof.
In the incidence diagram
[…]
the universal line $P_{F}$, as a correspondence in $A_{5}(F\times X)$, gives a map
$P_{*}:h(F)(1)\to h(X)$
By the results in [deC-M,Thm.6.2 ] $h(S)$ is a direct summand of $h(S^{[2]})\simeq h(F)$. Therefore we get a map
[…]
Let
$h(S)\simeq\mathbf{1}\oplus\mathbf{L}^{\oplus\rho(S)}\oplus t_{2}(S)\oplus\mathbf{L}^{2}$
be a refined Chow-Künneth decomposition, as in [KMP]. By composing with the inclusion $t_{2}(S)(1)\to h(S)(1)$ and the surjection $h(X)\to t(X)$ we get a map of motives in $\mathcal{M}_{{\operatorname{rat}}}(\mathbf{C})$,
$f:t_{2}(S)(1)\to t(X)$
For two distinct points $x,y\in S$ let us denote by $[x,y]\in S^{[2]}\simeq F$ the point of $F$ that corresponds to the subscheme $x\cup y\subset S$. If $x=y$ then $[x,x]$ denotes the element in $A^{4}(F)$ represented by any point corresponding to a non reduced subscheme of length 2 on $S$ supported on $x$. With these notations the special degree 1 cycle $c_{F}\in A^{4}(F)$ (see [SV,Lemma A3]), given by any point on a rational surface $W\subset F$, is represented by the point $[c_{S},c_{S}]\in F$, where $c_{S}$ is the Beauville-Voisin cycle in $A_{0}(S)$ such that $c_{2}(S)=24c_{S}$. We also have (see [SV,Prop. 15.6])
$(A_{0}(F)_{2})_{hom}=<[c_{S},x]-[c_{S},y]>,$
where $A_{0}(F)=A_{0}(F)_{0}\oplus A_{0}(F)_{2}\oplus A_{0}(F)_{4}$ is the Fourier decomposition, as in [SV].
We claim that the map $\phi:A_{0}(S)\to A_{0}(S^{[2]})=A_{0}(F)$ sending $[x]$ to $[c_{S},x]$ is injective and hence the surjective map $A_{0}(S)_{0}\to(A_{0}(F)_{2})_{hom}$, sending $[x]-[y]$ to $[c_{S},x]-[c_{S},y]$ gives an isomorphism
$A_{0}(S)_{0}\simeq(A^{4}(F)_{2})_{hom}.$
The variety $S^{[2]}$ is the blow-up of the symmetric product $S^{(2)}$ along the diagonal $\Delta\cong S$. Let $\tilde{S}$ be the inverse image of $\Delta$ in $S^{[2]}$. Then $\tilde{S}$ is the image of the closed embedding $s\to[c_{S},s]$. By a result proved in [Ba,Thm.2.1] the induced map of 0-cycles $A_{0}(\tilde{S})\to A_{0}(S^{[2]})$ is injective. Therefore the map $\phi$ is injective.
From the isomorphisms $A_{0}(S)_{0}\simeq(A_{0}(F)_{2})_{hom}$ and $A_{0}(F)_{2})_{hom}\simeq A_{1}(X)_{hom}$ (see [BP]) we get
$A^{3}(t_{2}(S)(1)=A^{2}(t_{2}(S))=A_{0}(S)_{0}\simeq A^{3}(X)_{hom}\simeq A^{3}(t(X))$
Since $A^{i}(t_{2}(S)(1))=A^{i}(t(X))=0$ for $i\neq 3$ the map $f:t_{2}(S)(1)\to t(X)$ gives an isomorphism on all Chow group, see [BP,Lemma 2.3]. Therefore $t_{2}(S)(1)\simeq t(X)$
Let $X\in\mathbf{P}^{5}_{\mathbf{C}}$ be a smooth cubic fourfold and let $F=F(X)$ be its Fano variety of lines. The irreducible holomorphic symplectic variety $F$ is deformation equivalent to the Hilbert scheme $S^{[2]}$ of two points on a K3 surface $S$. A finite order automorphism $\sigma$ of $S$ induces an automorphism $\sigma^{[2]}$ on the Hilbert scheme $S^{[2]}$ that is called natural.
Definition 5.5.
An automorphism $\sigma$ of $F(X)$ is called standard if $(F(X),\tilde{\sigma})$ is deformation equivalent to $(S^{[2]},\phi^{[2]})$, where $\phi$ is an automorphism of $S$.
Example 5.6.
(1) The Klein cubic fourfold $X$, whose equation is of the form
$x^{2}_{0}x_{1}+x^{2}_{1}x_{2}+x^{2}_{2}x_{3}+x^{2}_{3}x_{4}+x^{2}_{4}x_{0}+x^{3}_{5}=0,$
admits a symplectic automorphism of order 11 which cannot be standard because there is no symplectic automorphism of order 11 on a K3 surface.
(2) The non-symplectic automorphisms of order 3 on the cubic fourfolds belonging to the families $V_{1}$ and $V_{3}$ in Sect.3 are non-standard, see [BCS 1,Cor.7.6]
Definition 5.7.
Let $X\in\mathbf{P}^{5}_{\mathbf{C}}$ be a smooth cubic fourfold and let $F(X)$ be its Fano variety of lines equipped with the polarization $g$ which is by definition the class in $H^{2}(F,\mathbf{Z})$ of the restriction to $F$ of the Plucker line bundle $\mathcal{L}$ on the Grassmannian $G(2,6)$ of lines in $\mathbf{P}^{5}$. An automorphism $\psi$ of $F(X)$ is polarized if it preserves the Plucker polarization, i.e. $\psi^{*}g=g$.
An automorphism of $F(X)$ is polarized if and only if it is induced from an automorphism of the cubic fourfold $X$, see [Fu,Lemma1.2].
Lemma 5.8.
Let $(S,L)$ be a polarized K3 surface of genus $g$ and degree $d$, with $L^{2}=2g-2=d$ and $d=2(n^{2}+n+1)$. Let $X\in\mathcal{C}_{d}$ be the corresponding cubic fourfold in 5.1. Let $\sigma$ be a symplectic automorphism of $S$ of prime order $p$ and let $\sigma^{[2]}$ be the induced automorphism on $S^{[2]}\simeq F(X)$. The natural automorphism $\sigma^{[2]}$ on $F(X)$ is polarized and hence it is induced by an automorphism $\tau$ on $X$. The fixed locus of the automorphism $\tau$ on $X$ consists of a finite number of isolated points.
Proof.
If $\sigma$ is a symplectic automorphism of $S$ of prime order $p$, then $p=3,5,7$. There is an isomorphism
$H^{2}(F(X),\mathbf{Z})\simeq H^{2}(S^{[2]},\mathbf{Z})\simeq H^{2}(S,\mathbf{Z})\oplus\mathbf{Z}\delta,$
Here $\delta=(1/2)|\Delta|\in H^{2}(S^{[2]},\mathbf{Z})$, with $\Delta\subset S^{[2]}$ the divisor consisting of zero- dimensional closed subschemes supported only at a single point. Then
(5.9) $\Delta=\mathbf{P}(T_{S})=\bigcup_{p\in S}\Delta_{p},$
where $\mathbf{P}(T_{S})$ is the geometric projectivization of the tangent bundle $T_{S}$. For each $p\in S$ $\Delta_{p}$ is the rational curve consisting of 0-dimensional subschemes $\xi\in\Delta$ such that $\operatorname{supp}(\xi)=p$, see [FV,2.1].
The Fano variety $F(X)$ has a polarization $g$, where
$g=2f-(2n+1)\delta,$
with $f=|L|\in H^{2}(S,\mathbf{Z})$, see [Hass, Thm.29]. The symplectic automorphism $\sigma$ of the polarized K3 surface $S$ fixes the classes $f$ and the automorphism $\sigma^{[2]}$ of $S^{[2]}$ fixes the class of $\Delta$. Therefore the induced automorphism fixes the polarization $g$. Therefore the natural automorphism $\sigma^{[2]}$ on $F(X)$ is polarized and hence it is induced by an automorphism $\tau$ of $X$.
The automorphism $\sigma^{[2]}$ has a finite number $m_{p}$ of fixed points, with $m_{3}=27,m_{5}=14,m_{7}=9$, see [Bois, Ex.2]. Every subspace $\xi\in S^{[2]}$ fixed under $\sigma^{[2]}$ is the union of reduced subspaces with disjoint supports of the form
(1) $\{p\}$ with $\sigma(p)=p$ ,
(2)$\xi_{p}$ with $\xi_{p}$ a non reduced subscheme of length 2 supported on $p$ with $\sigma(p)=p$, see [Bois, 4.2 and Rk. 6].
The fixed locus of $\sigma$ on $S$ consists on a finite number of isolated points and and the fixed 1-dimensional subschemes $\xi_{p}$ in (2) correspond to $\Delta_{p}\subset S^{[2]}$, with $\sigma(p)=p$.
Let $x\in X$ and let $\tau$ be the automorphism induced on $X$. Let $\tau(x)=x$ and let $l\in F(X)$ be a line trough $x$. Then also the line $\sigma^{[2]}(l)$ passes trough $x$. Therefore the 1-dimensional subscheme $C_{x}\subset F(X)$ of all lines trough $x$ corresponds, under the isomorphism $F(X)\simeq S^{[2]}$, to a 1-dimensional subscheme fixed under $\sigma^{[2]}$. So there are only a finite number of points on $X$ fixed under $\tau$. ∎
5.1. The case p=3
[…]
the surface $\tilde{S}$ is the blow-up of $S$ at the six fixed points $P_{i}$ of $\sigma$.
The family $\mathcal{S}$ of polarized K3 surface with a symplectic automorphism of order 3 is the union of countably many components of dimension 7. Similarly the family $\mathcal{T}$ of K3 surfaces that are quotients of K3 surfaces with a symplectic automorphism of order 3 is the union of countably many components of dimension 7. The correspondence between K3 surfaces in $\mathcal{S}$ and in $\mathcal{T}$ has been described in [GM]. A K3 surface $S$ is a generic member of the component $\mathcal{S}_{2d}\subset\mathcal{S}$ of those surfaces having a polarization $L$ of degree $2d$ if and only if the corresponding K3 surface $Y$ is a generic member of the component $\mathcal{T}_{6d}\subset\mathcal{T}$ of those surfaces having a polarization $H$ of degree $6d$. Conversely a K3 surface $Y\in\mathcal{T}_{2e}\subset\mathcal{T}$ with a polarization of degree $2e$ corresponds to a K3 surface $S\in\mathcal{S}_{6e}\subset\mathcal{S}$ with a polarization of degree $6e$, see see [GM,Cor.5.3].
Let $S\in\mathcal{F}_{g}$, where $g=n^{2}+n+2$ with a symplectic automorphism $\sigma$ of order 3 and let $X\in\mathcal{C}_{d}$ be the image of $S$ under the rational map in 5.1. Then $F(X)\simeq S^{[2]}$ and $\sigma$ induces a natural automorphism $\sigma^{[2]}$ on $F(X)$ with 27 isolated fixed points. The automorphism $\sigma^{[2]}$ is induced by an automorphism $\tau$ of $X$, which in turn is induced by a linear automorphism of $\mathbf{P}^{5}$. Then one can choose coordinates $[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]$ in such a way that $(X,\tau)$ belongs to one of the families of cubic fourfolds with a symplectic automorphism of order 3, as described in [Fu, Thm.1.1]. There are two families of cubic fourfolds with a symplectic automorphism of order 3, whose fixed locus on $F(X)$ consists of 27 isolated fixed points.
The first family $F_{3}$ has dimension 8. Every cubic in $F_{3}$ is invariant under the automorphism of $\mathbf{P}^{5}$
(5.11) $\tau_{1}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to[x_{0},x_{1},x_{2},x_{3},\zeta x_{4},\zeta^{2}x_{5}]$
with $\zeta^{3}=1$ and has an equation of the form
$F(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})=f(x_{0},x_{1},x_{2},x_{3})+x^{3}_{4}+x^{3}_{5}+x_{4}x_{5}L(x_{0},x_{1},x_{2},x_{3})=0,$
with $f$ of degree 3 and $L$ of degree 1. The fixed locus of the automorphism $\tau_{1}$ on $X$ is the cubic surface $S:f(x_{0},x_{1},x_{2},x_{3})=0$. Since this fixed locus is not finite the automorphism induced by 5.11 on $F(X)$ cannot coincide with $\sigma^{[2]}$ of $F(X)$, hence it is not natural. In [Mong,7.2.8] it is proved that the order 3 symplectic automorphism of the cubics in $F_{3}$ are standard.
The second family $G_{3}$ has dimension 8 . Every cubic in $G_{3}$ is invariant under the automorphism of $\mathbf{P}^{5}$
$\tau_{2}:[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]\to[x_{0},x_{1},\zeta x_{2},\zeta x_{3},\zeta^{2}x_{4},\zeta^{2}x_{5}]$
and has an equation of the form
$F(x_{0},x_{1},x_{2},x_{3},x_{4},x_{5})=f_{1}(x_{0},x_{1})+f_{2}(x_{2},x_{3})+f_{3}(x_{4},x_{5})+\sum_{i,j,k}a_{ijk}x_{i}x_{j}x_{k}$
where $f_{1},f_{2},f_{3}$ have degree 3 and $i=0,1;j=2,3;k=4,5$. The fixed locus of $\tau_{2}$ on $X$ is finite and consists of the 9 isolated points $P_{i,1},P_{i,2},P_{i,3}$ satisfying the the equation $f_{i}=0$, for $i=1,2,3$ The fixed locus on $F(X)$ corresponds to the 27 lines $\overline{P_{i,k}P_{j,l}}$, for $1\leq i<j\leq 3$ and $k,l=1,2,3$, see [Fu,IV-(5)]. Each point on $X$ fixed by the automorphism is the intersection of 3 lines in $F(X)$ fixed by the automorphism.
Since the moduli space of polarized K3 surfaces with a symplectic automorphism of order 3 has dimension 7 (see [GS],Cor.5.1]), the image in $G_{3}$ has codimension 1 and contains the cubics $X$ such that the automorphism $\tau_{2}$ is natural.
Let $S$ be K3 surface with a polarization $L$ of degree $L^{2}=2d=2g-2$, where $d=n^{2}+n+1$, with a symplectic automorphism $\sigma$ of order 3. Then a desingularization $Y$ of $S/\sigma$ has a polarization $H$ of degree $6d$ and genus $g^{\prime}=3n^{2}+3n+4$. Similarly if $S$ is a K3 surface with a polarization $L$ of degree $L^{2}=6e=2g-2$, where $g=m^{2}+m+2$, then the desingularization $Y$ of $S/\sigma$ has a polarization $H$ of degree $2e$ and genus $g^{\prime}=(m^{2}+m+1)/3+1$. For $m=4$ we get $g=22$ and $g^{\prime}=8$. Therefore $S\in\mathcal{F}_{22}$ and $Y\in\mathcal{F}_{8}$
Example 5.12.
Let $S$ be a polarized K3 surface in $\mathcal{F}_{22}$, with a symplectic automorphism $\sigma$ of order 3 and let $\tilde{X}\in\mathcal{C}_{42}$ be the cubic fourfold corresponding to $S$ under the map in 5.1. There is an isomorphism $F(\tilde{X})\simeq S^{[2]}$ and an order 3 automorphism $\tau$ of $\tilde{X}$ induced by the automorphism $\sigma^{[2]}$ of $F(\tilde{X})$. Let $Y$ be a desingularization of $S/\sigma$ and let $X\in\mathcal{C}_{14}$ be the cubic fourfold corresponding to $Y$ under the map $\mathcal{F}_{8}\dashrightarrow\mathcal{C}_{14}$.Then
(5.13) $F(\tilde{X})\simeq S^{[2]}\ ,\ F(X)\simeq Y^{[2]}.$
The 8-dimensional family $G_{3}$ of cubic fourfolds with a symplectic automorphism of order 3 contains two divisors , one corresponding to the cubics $\tilde{X}$ in $\mathcal{C}_{42}$ and one to the Pfaffian cubics $X\in\mathcal{C}_{14}$.
Since $42\equiv 0(6)$ if $S$ and $S^{\prime}$ have the same image in $\mathcal{C}_{42}$, under the rational map $\mathcal{F}_{22}\dashrightarrow\mathbf{C}_{42}$, then $S$ and $S^{\prime}$ are Fourier- Mukai partners, see [Brak,Thm.1.1]. Therefore $D^{b}(S)\simeq D^{b}(S^{\prime})$ and the Chow motives $h(S)$ and $h(S^{\prime})$ are isomorphic.
The quotient map $S\to S/\sigma$ and the resolution $Y\to\ S/\sigma$ give a generically 3:1 rational map $\pi:S\dashrightarrow Y$ and therefore a rational map
$\pi^{[2]}:S^{[2]}\dashrightarrow Y^{[2]},$
which is defined outside the 27 fixed points on $S^{[2]}$. The rational map $\pi^{[2]}$ induces, via the isomorphisms in 5.13, a rational map $F(\tilde{X})\dashrightarrow F(X)$.
Since the transcendental motive $t_{2}(-)$ of a surface is a birational invariant the maps $\tilde{S}\to S$ and $\tilde{S}\to Y$ in 5.10 induce a map $\theta:t_{2}(S)\to t_{2}(Y)$ that is the projection onto a direct summand, see [Ped,Prop.1]. Therefore
$t_{2}(S)=t_{2}(Y)\oplus N.$
By a result of Huybrechts a symplectic automorphism acts trivially on the Chow group of 0-cycles $A_{0}(S)=CH_{0}(S)\otimes\mathbf{Q}$. Since $A_{0}(t_{2}(S)=A_{0}(S)_{hom}$ , $A_{i}(t_{2}(S))=0$, for $i\neq 0$, and $A_{0}(S)^{\sigma^{*}}=A_{0}(S)$ we get $A_{i}(N)=0$, for all $i$. Therefore $N=0$,and hence
$t_{2}(S)=t_{2}(Y)$
The cubic fourfolds $\tilde{X}$ and $X$ are both rational.The transcendental motives $t(\tilde{X})$ and $t(X)$ are isomorphic because $t_{2}(S)(1)\simeq t(\tilde{X})$ and $t_{2}(Y)(1)\simeq t(X)$ (see 5.3), with $t_{2}(S)\simeq t_{2}(Y)$ Therefore
(5.14) $t_{2}(S)(1)\simeq t(\tilde{X})\simeq t(X).$
The K3 surface $S$ is associated to the motive of $X$ and of $\tilde{X}$ in the sense of 1.9.
Let $\mathcal{A}_{\tilde{X}}$ and $\mathcal{A}_{X}$ be the Kuznetsov components of $\tilde{X}$ and $X$, respectively. Since $X$ and $\tilde{X}$ belong to a Hassett divisors $\mathcal{C}_{d}$, with d satisfying the numerical condition in (*), there are isomorphisms
$\mathcal{A}_{\tilde{X}}\simeq D^{b}(S)\ ;\ \mathcal{A}_{X}\simeq D^{b}(Y),$
where
$D^{b}(\tilde{X})=<\mathcal{A}_{\tilde{X}},\mathcal{O}_{\tilde{X}},\mathcal{O}_{\tilde{X}}(1),\mathcal{O}_{\tilde{X}}(2)>.$
and similarly for a semi-orthogonal decomposition of $D^{b}(X)$. Let $G=<\sigma>$ and let $D^{b}_{G}(S)$ be the derived category of $G$-equivariant coherent sheaves on $S$. Then
$D^{b}(Y)\simeq D^{b}_{G}(S),$
see [XH, 3.1].